Write and or solve the system of equations using any method.
.2x-.5y=-3.8
.3x+.4y=10.4
Thanks in advance!

Answer by mananth(16949)

0.20 x -0.50 y = -3.80 .............1
0.30 x + 0.40 y = 10.40 .............2
Eliminate y
multiply (1)by 0.4
Multiply (2) by 0.5
0.08 x -0.20 y = -1.52
0.15 x + 0.2 y = 5.20
Add the two equations
0.23 x = 3.68
/ 0.23
x = 16
plug value of x in (1)
0.2 x -0.5 y = -3.8
3.2 -0.5 y = -3.8
-0.5 y = -3.8 -3.2
-0.5 y = -7
y = 14